6.74g of the monoprotic acid KHP(MW=204.1g/mol) is dissolved into water. The sample is titrated with a 0.703 M solution of calcium hydroxide to the equivalence point. What volume of base was used?


1
Expert's answer
2021-10-28T02:57:40-0400

2KHP + Ca(OH)2 → Ca(KP)2 + 2H2O

n(KHP) "= \frac{6.74}{204.1} = 0.033 \;mol"

n(Ca(OH)2) = "\frac{1}{2}" n(KPH) "= \frac{1}{2} \times 0.033 = 0.016 \;mol"

Proportion:

0.703 mol – 1000 mL

0.016 mol – x mL

"x = \frac{0.016 \times 1000}{0.703} = 23.48 \;mL"

Answer: 23.48 mL